# MCHex: Marching Cubes Based Adaptive Hexahedral Mesh Generation with Guaranteed Positive Jacobian

### v2

@@ 7 Conclusion and Future Work
Some limitations and future directions are also identified. First, the midpoint subdivision strategy inherently increases hex element count, imposing greater computational burden in subsequent numerical simulations. This limitation could be mitigated by adopting the vertex-efficient refinement approach in Figure 14, which prevents the midpoint subdivision from propagating to all internal cells. Second, the current theoretical guarantee of positive Jacobian is restricted to basic 3-refinement templates, which can lead to excessive refinement. Extending this proof to improved 3-refinement or 2-refinement templates would substantially reduce element counts while maintaining quality guarantees. Third, the method cannot preserve sharp features while ensuring positive Jacobian, a limitation inherent to the standard MC algorithm. A recent work demonstrates that incorporating additional feature-aware vertices on the isosurface and combining with deep learning can effectively preserve sharp edges (chen2021neural). Adapting the proposed tessellation templates to redesign templates in Figure 6 and Figure 7 would possibly enable sharp feature preservation within this framework.
@@ Appendix A Positive Jacobian Check
Due to the rotational symmetry of the two polyhedra in Case 12 with respect to a $180^{\circ}$ rotation around the z-axis, it is unnecessary to recompute the offsets for the polyhedron on the opposite side. Instead, the corresponding offsets can be obtained by rotating the cell by $180^{\circ}$ around the z-axis, looking up the precomputed offsets $\mathbf{v}=(a,b,c)$, and applying $\mathbf{v}^{\prime}=(-a,-b,c)$ as the new offsets. Additionally, Case 12 has a mirror symmetry along the x-direction, a case not shown in Figure 6 but illustrated as Case 11 of Figure 1 in (chernyaev1995marching). Under this mirror transformation, where vertex coordinates are mapped $(x,y,z)\to(-x,y,z)$, the corresponding offset becomes $\mathbf{v}^{\prime\prime}=(-a,b,c)$.
@@
Finally, two additional observations are noted regarding this proof. First, the conclusion that “midpoint subdivision guarantees a positive Jacobian” for Cases 1 to 11 and Case 13 can be extended to the entire parameter space $\xi,\eta,\zeta\in[0,1]^{3}$, not only at the eight corner points and the center point. However, extending this proof to Case 12 and Case 14 is challenging. The piecewise offset approach described above has so far failed to establish a full parametric proof. If such a proof can be achieved in the future, it would rigorously ensure that the generated hex elements can be directly applied in engineering simulations. Second, beyond the transition cells shown in Figure 8, a limited number of transition cells from other methods (ito2009octree, owen2017template, livesu2021optimal) are also examined. However, a complete proof is not feasible due to the high complexity of configurations. The bitter lesson learned here is that the closer the mapping from the unit cube to the transition cell is to an affine transformation, such as when all four points of each face of the transition cell are coplanar, the more tractable the proof becomes. This is why, in the first paragraph of Section 5, it is suggested that the transition cells from (ito2009octree, owen2017template, zhang2013robust) are the most promising candidates for future extensions. For more advanced methods such as (livesu2021optimal), it may be necessary to devise more sophisticated midpoint subdivision definitions to enable positive Jacobian guarantees.

### v3

@@ 7 Conclusion and Future Work
Some limitations and future directions are also identified. First, the midpoint subdivision strategy inherently increases hex element count, imposing greater computational burden in subsequent numerical simulations. This limitation could be mitigated by adopting the vertex-efficient refinement approach in Figure 14, which prevents the midpoint subdivision from propagating to all internal cells. Second, the current theoretical guarantee of positive Jacobian is restricted to basic 3-refinement templates, which can lead to excessive refinement. Extending this proof to improved 3-refinement or 2-refinement templates would substantially reduce element counts while maintaining quality guarantees. Third, the method cannot preserve sharp features while ensuring positive Jacobian, a limitation inherent to the standard MC algorithm. A recent work demonstrates that incorporating additional feature-aware vertices on the isosurface and combining with deep learning can effectively preserve sharp edges (8). Adapting the proposed tessellation templates to redesign templates in Figure 6 and Figure 7 would possibly enable sharp feature preservation within this framework.
@@ Appendix A Positive Jacobian Check
Due to the rotational symmetry of the two polyhedra in Case 12 with respect to a $180^{\circ}$ rotation around the z-axis, it is unnecessary to recompute the offsets for the polyhedron on the opposite side. Instead, the corresponding offsets can be obtained by rotating the cell by $180^{\circ}$ around the z-axis, looking up the precomputed offsets $\mathbf{v}=(a,b,c)$, and applying $\mathbf{v}^{\prime}=(-a,-b,c)$ as the new offsets. Additionally, Case 12 has a mirror symmetry along the x-direction, a case not shown in Figure 6 but illustrated as Case 11 of Figure 1 in (9). Under this mirror transformation, where vertex coordinates are mapped $(x,y,z)\to(-x,y,z)$, the corresponding offset becomes $\mathbf{v}^{\prime\prime}=(-a,b,c)$.
@@
Finally, two additional observations are noted regarding this proof. First, the conclusion that “midpoint subdivision guarantees a positive Jacobian” for Cases 1 to 11 and Case 13 can be extended to the entire parameter space $\xi,\eta,\zeta\in[0,1]^{3}$, not only at the eight corner points and the center point. However, extending this proof to Case 12 and Case 14 is challenging. The piecewise offset approach described above has so far failed to establish a full parametric proof. If such a proof can be achieved in the future, it would rigorously ensure that the generated hex elements can be directly applied in engineering simulations. Second, beyond the transition cells shown in Figure 8, a limited number of transition cells from other methods (21, 33, 27) are also examined. However, a complete proof is not feasible due to the high complexity of configurations. The bitter lesson learned here is that the closer the mapping from the unit cube to the transition cell is to an affine transformation, such as when all four points of each face of the transition cell are coplanar, the more tractable the proof becomes. This is why, in the first paragraph of Section 5, it is suggested that the transition cells from (21, 33, 51) are the most promising candidates for future extensions. For more advanced methods such as (27), it may be necessary to devise more sophisticated midpoint subdivision definitions to enable positive Jacobian guarantees.